There are 72 members of the show choir. There are 6 more boys than girls in the choir. How many girls and boys are in the choir?
Found 2 solutions by fractalier, ankor@dixie-net.com:
Answer by fractalier(6550)

We can write
b + g = 72 and
b = g + 6
Substitute the second into the first and get
g + 6 + g = 72
2g + 6 = 72
2g = 66
g = 33 girls
b = 39 boys

Answer by ankor@dixie-net.com(22740)

b = no. of boys
g = no. of girls
:
There are 72 members of the show choir.
b + g = 72
:
There are 6 more boys than girls in the choir.
b = g+6
:
How many girls and boys are in the choir?
replace b with (g+6) in the 1st equation
(g+6) + g = 72
subtract 6 from both sides
2g = 72 - 6
g = 66/2
g = 33 girls
:
:
To confirm our answer, find the no. of boys
b = 33 + 6
b = 39 boys
then
39 + 33 = 72